Let๐›ผ(๐‘ฅ)=๐‘ฅ and define๐‘“ as๐‘“(๐‘ฅ)=1 if๐‘ฅ is rational and๐‘“(๐‘ฅ)=0 if๐‘ฅ is irrational. Find ๐ผ(๐‘“, ๐›ผ) and ๐ผ(๐‘“, ๐›ผ).ย 


1
Expert's answer
2022-02-03T16:09:04-0500

By definition;

"\displaystyle
\alpha(x)=x \text{ and }\\
f:[a,b]\rightarrow\R \text{ is defined as }
f(x)=\begin{cases}1&\text{if }x\text{ is rational}\\0&\text{if }x\text{ is irrational}\end{cases}"

Now, for every partition P of "\displaystyle
[a,b],"

"\displaystyle
M_k(f) = \sup \{f(x) : x \in [x_{k-1}, x_k]\}=1\\
m_k(f) = \inf \{f(x) : x\in[x_{k-1}, x_k]\}=0", since every subinterval contains both rational and irrational numbers.


Thus,

"\displaystyle
U(P, f, a) =\sum^n_{k=1} M_k(f) \Delta \alpha_k=1\\\text{and}"

"\displaystyle
L(P, f, \alpha) =\sum^n_{k=1} m_k(f) \Delta \alpha_k=0\ \ \forall\ \text{partition }P.\\
\text{Where }\Delta\alpha_k=\Delta\alpha(x_{k})=\alpha_k-\alpha_{k-1}."


Hence, it follows that we have;

"\displaystyle
\underline{I}(f,\alpha)=\underline{\int}^b_af\ dx=0\\
\quad\\
\overline{I}(f,\alpha)=\overline{\int}^b_af\ dx=1"